Find an equation for the line that goes through the point (5,4) and is perpendicular to the line satisfying -5x-5y=-47
what is the actual formula for y as a function of x?

Answer by ewatrrr(24785)


Hi
the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
for ex: -5x - 5y = -47: | y = -x + 47/5
Slope m = -(1/1) LIne perpendicular to it wil have a slope m = 1 (negative reciprocal)
New Line:
y = 1*x + b |using ordered pair Pt(5,4) to solve for b
4 = 5 + b
-1 = b
y = x -1